OpenStudy (gabylovesyou):

What are the approximate solutions of 2x2 + 9x = 8 to the nearest hundredth? x ≈ 1.22 and x ≈ 3.28 x ≈ −0.76 and x ≈ 5.26 x ≈ −1.22 and x ≈ −3.28 x ≈ −5.26 and x ≈ 0.76

OpenStudy (gabylovesyou):

@phi

OpenStudy (phi):

Can you write this in standard form \[ 2x^2 + 9x = 8 \]? standard form is \( a x^2 + b x + c =0 \)

OpenStudy (phi):

add -8 to both sides and simplify. what do you get ?

OpenStudy (gabylovesyou):

2x^2 + 9x - 8 = 8 - 8 x(2x + 9) = 8

OpenStudy (phi):

1st line is ok. but you should remember what we are trying to do: make it look like a x^2 + bx + c =0 2x^2 + 9x -8 = 8-8 on the right side, 8-8 is 0 , so you can simplify to get 2x^2 + 9x -8 =0 that is standard form. the only thing you had to do is replace 8-8 with 0.

OpenStudy (phi):

now the hard part. To find the solutions you must use the quadratic formula do you know it ?

OpenStudy (gabylovesyou):

no i dont ;/

OpenStudy (phi):

it is complicated: \[ x = \frac{-b ± \sqrt{b^2 - 4 ac}}{2a} \] to use it you must identify the numbers that go with a, b , and c match your equation 2x^2 + 9x + -8 =0 with ax^2 + bx + c = 0 what is a, b and c ?

OpenStudy (gabylovesyou):

2, 9 and -8

OpenStudy (gabylovesyou):

@Mertsj

OpenStudy (phi):

yes, now use those numbers to replace the letters in the formula. can you do that ?

OpenStudy (phi):

wherever you see an a, put in a 2 instead. replace b with 9 and c with -8

OpenStudy (gabylovesyou):

but how the heck do i solve that LOL xD

OpenStudy (phi):

The longest journey begins with a single step... The first step is replace a with 2. I will do it to get you started \[ x = \frac{-b ± \sqrt{b^2 - 4 ac}}{2a} \] becomes \[ x = \frac{-b ± \sqrt{b^2 - 4\cdot 2\cdot c}}{2 \cdot 2} \] can you replace b with 9 ? don't worry about solving anything. Just replace the letters with numbers.

OpenStudy (gabylovesyou):

NEVERMIND I GOT A CALCULATOR lol is it D?

OpenStudy (gabylovesyou):

@phi

OpenStudy (phi):

to finish this problem. replace b with 9 in \[ x = \frac{-b ± \sqrt{b^2 - 4\cdot 2\cdot c}}{2 \cdot 2} \] you get \[ x = \frac{-9 ± \sqrt{9^2 - 4\cdot 2\cdot c}}{2 \cdot 2} \] now replace c with -8 to get \[x = \frac{-9 ± \sqrt{9^2 - 4\cdot 2\cdot -8}}{2 \cdot 2}\] you need to know your order of operations to simplify. you will get \[ x= \frac{-9± \sqrt{145}}{4} \] the two roots will be \[ x= \frac{-9+ \sqrt{145}}{4} \] and \[ x= \frac{-9 - \sqrt{145}}{4} \] using a calculator we will find x = (-9 + 12.0416)/4= 0.7604 and x= (-9 - 12.0416)/4 = -5.26 choice D